# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ 3.3 Shifting regularization from $H$ to $g$
Figure 2: The influence of the penalty parameter $\lambda$ in (15), where we have replaced $y$ with $1/x$.
@@
$\{(g,\sigma_{y},u,y)\mid Hg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),{H}\in\hat{\mathcal{C}}_{\mathrm{opt}}\}$
@@
$\{(g,\sigma_{y},u,y)\mid Dg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),h(g)=0\},$
where $D$ is a fixed trajectory matrix. We next move the constraint on $g$ to the cost via regularization, leading to a direct DDPC
@@
$\mathrm{subject~to}$ $Dg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
@@
Similar to (13), one may require the regularization parameter $\lambda_{w}$ in (16) to approach infinity. Fortunately, in our context, we can establish an exact penalty with a finite regularization parameter $\lambda_{w}$. In particular, (16) is a quadratic optimization problem of the form
@@
Consider (17) and (18), where $\mathcal{X}$ is a convex set, $M$ is positive semidefinite, and $x_{1}^{{\mathsf{T}}}Mx_{1}$ is Lipschitz continuous with Lipschitz constant $L$ with respect to $\|\cdot\|_{p}$ over $\mathcal{X}$. If there exists an optimal solution $(x_{1}^{*},x_{2}^{*})$ for (17) such that $x_{1}^{*}$ is an interior point of $\mathcal{X}$, then there exists $\lambda_{w}^{*}>0$, such that (17) and (18) are equivalent for all $\lambda_{w}>\lambda_{w}^{*}$, i.e., they have same optimal solutions.
One key fact in this theorem is that we can find a finite penalty parameter such that (17) and (18) are equivalent. Theorem 1 confirms that $\|Dx_{2}\|_{p}$ is an exact penalty function for the quadratic problem (17). To our knowledge, this result is new and might be of independent interest. The key proof idea is to show that the optimal solution of (17) is also a local minimizer of (18). Since (18) is a convex problem, any local minimizer is also a global minimizer. Consequently, $\lambda_{w}\|Dx\|_{p}$ serves as an exact penalty term for some finite weight $\lambda_{w}>\lambda_{w}^{\star}$. The proof details are technically involved, and we present the full proof in Appendix A.
To illustrate Theorem 1, we randomly generate problem instances of (17) and (18). As shown in Figure 3, all $p$-norms render $\|Dx\|_{p}$ an exact penalty. However, smaller values of $p$ lead to faster convergence of the optimal value of (18) toward that of (17) as $\lambda_{w}$ increases. Since $\|Dx\|_{p}$ decreases monotonically with $p$, larger $p$ values impose weaker penalties and therefore require larger $\lambda_{w}$ to guarantee equivalence between the two problems.
@@ 4 Regularization in direct data-driven control via convex relaxations
In this section, we apply Theorem 1 and derive three direct DDPC strategies via convex relaxations, including regularization using projection-based norm, causality-based norm, and $l_{1}$ norm. The projection-based norm and $l_{1}$ norm are widely used in the literature (see [1] and its references), and the causality-based norm is new.
@@ 4.1 Projection-based norm and a Hankel-form SPC
which only considers the SPC constraint (10b) for the inner identification. This bi-level DDPC (19) is much simpler than the general one (7). However, similar to Example 1, the optimal value of (19) may not be lower than that of (7); in other words, (19) is not a relaxation of (7).
@@
$\Pi_{1}:=H_{1}^{\dagger}H_{1},\qquad H_{1}:=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}).$ (20)
@@
Note that (21) differs slightly from the classical SPC formulation and we call it Hankel-form SPC. In particular, the classical SPC does not involve the use of a Hankel matrix; instead, it employs an explicit multi-step predictor (also see Remark 2). For self-completeness, we provide further details in Appendix B.1.
Following the discussion in Section 3.3, we here fix $H=H_{\mathrm{d}}$ and impose a constraint on $g$ to get an equivalent reformulation for (21). In this case, we use a projection-based norm $\|(I-\Pi_{1})g\|_{2}$ as in [1], and consider
@@
The proof is provided in Appendix B.2. The key idea is to show that feasible regions for decision variables $u,y,\sigma_{y}$ are the same for (19), (21) and (22). Note that the data matrix $H_{\textnormal{d}}$ in Proposition 1 can be arbitrary and may come from a nonlinear system.
It is now clear that (22) is in the form of (17). Thus, applying Theorem 1 directly leads to the following result.
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},$ and $\mathcal{Y}$ are compact and convex sets. If the optimal solution $(\sigma_{y}^{\star},u^{\star},y^{\star})$ to (22) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$ such that
@@
is equivalent to (22) for all $\lambda_{g}>\lambda_{g}^{*}$. In other words, they have the same optimal cost value and optimal solution.
We provide some details in Appendix B.4. We call (23) as Linearity-based DDPC (L-DDPC). Combining Corollary 1 and Proposition 1 gives us the following result.
@@
Consider the bi-level DDPC (19) and the single-level L-DDPC (23). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are compact and convex sets and the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (19) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$. Then, for any nonnegative $\lambda_{g}\geq 0$, (23) is a convex relaxation of (19), that is:
@@
Our Theorem 2 is closely related to [1, Theorem 4.6], but there are two key differences. Conceptually, as illustrated in Example 1, the L-DDPC (23) is only a convex relaxation of (19), and not necessarily of the original bi-level problem (7). This subtle distinction was overlooked in [1, Theorem 4.6]. Technically, we establish that (23) serves as a convex relaxation of (19) for any nonnegative parameter $\lambda_{g}\geq 0$, rather than only for sufficiently small $\lambda_{g}$ as in [1, Theorem 4.6]. The threshold $\lambda_{g}^{*}$ that makes (19) and (23) equivalent depends on both the Lipschitz constant of the objective function and problem data (e.g., $H_{\textnormal{d}}$) rather than only on the Lipschitz constant as in [1, Theorem 4.6]. Moreover, our proof is elementary, relying on Theorem 1 and Proposition 1, which are inspired by the theoretical condition of partial calmness. We avoid utilizing [43, Proposition 2.4.3] as in [1, Theorem4.6], whose application in the DDPC setting may require additional adaptations.
@@ 4.2 Causality-based norm and a Causal SPC
which considers the SPC constraint (10b) and the causality constraint (10c) for the inner identification. If $H_{\textnormal{d}}$ has full row rank, the optimal solution $H^{*}_{\textnormal{sc}}$ for the inner problem in (24) is unique [31, Lemma 2], which can be obtained via its optimal solution $K^{*}_{\textnormal{sc}}$, derived as
$K_{\textnormal{sc}}^{*}=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1},$
@@
$H_{\textnormal{d}}=\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}\\ L_{31}&L_{32}&L_{33}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\end{bmatrix}$
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are convex sets. Fix any data matrix $H_{\textnormal{d}}$ in (5b) with full row rank and sufficiently large column number. The bi-level problem (24) has the same optimal value as the single-level problems (26) and (28). Furthermore, all three problems have the same unique optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$.
The proof is provided in Appendix B.3. It is clear that (28) is in the form of (17). Thus, applying Theorem 1 leads to the following result.
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U}$, and $\mathcal{Y}$ are compact and convex sets. If the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (28) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$, such that
@@
with $\hat{H}$ defined in (27), has the same optimal value and optimal solutions as (28) for all $\lambda_{g}>\lambda_{g}^{*}$.
@@
Consider the bi-level DDPC (24) and the single-level C-DDPC (29). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $H_{\textnormal{d}}$ has full row rank with sufficiently large column number. Let $\Gamma,\mathcal{U}$ and $\mathcal{Y}$ are compact and convex sets and the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (24) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$. For any positive $\lambda_{g}$, (29) is a convex relaxation of (24), that is:
@@
To our best knowledge, the single-level C-DDPC in (29) is new, and its relationship with the bi-level DDPC (24) has not been discussed before. The closest work on causality-based regularizer is [31], which regularizes on $\gamma$ in the $\gamma$-DDPC framework. Instead, our method (29) directly regularizes $g$, which can be more convenient to combine with $\|g\|_{1}$ for implicit trajectory selection.
@@ 4.3 Low-rank approximation and $l_{1}$ Norm
$H_{\textnormal{lr}}^{*}=\sum_{i=1}^{mL+n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$
@@
$\mathrm{subject~to}$ $\tilde{H}_{\textnormal{lr}}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
@@
The proof is immediate from the fact that the optimal solution for (31) is feasible for (32) for sufficiently large $\alpha_{\textnormal{c}}$ as $\mathrm{range}(H^{*}_{\textnormal{lr}})\subseteq\mathrm{range}(H_{\textnormal{d}})$.
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$, $\Gamma\times\mathcal{U}\times\mathcal{Y}$ is a convex set and the convex optimization problem (32) satisfies the Slater’s condition. Let $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ be an optimal solution for (32). There exists $\lambda_{g}^{*}\geq 0$, such that $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ is also an optimal solution for
@@
The proof is given in Appendix B.5. Propositions 3 and 4 imply that (33) is a relaxation of (30) with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ under mild conditions.
@@
Consider the bi-level problem (30) and the single-level problem (33). Suppose $Q\succ 0,R\succ 0$ and $\lambda_{y}>0$. Let $\Gamma,\mathcal{U}$ and $\mathcal{Y}$ be convex sets. There exists $\lambda_{g}^{*}\geq 0$, such that (33) is a convex relaxation of (30) with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$, that is:
@@
    any feasible $\sigma_{y},u,y$ for (30) is feasible to (33);
@@
    The value of $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ evaluated at the optimal solution of (33) is smaller than or equal to that of (30).
A similar statement as Theorem 4 appears in [1, Theorem 4.8]. However, we emphasize that the relaxation is only true with respect to the function $\|y\|_{Q}+\|u\|_{R}+\lambda_{y}\|\sigma_{y}\|$, which is treated differently in [1]. The reason is that while the optimal solution of (33) is the same as (32), their objective functions are different, which generally leads to different optimal values.

### v2

@@ 3.3 Shifting regularization from $H$ to $g$
Figure 2: Influence of the penalty parameter $\lambda$ in (15), where we have replaced $y$ with $1/x$.
@@
$\{(g,\sigma_{y},u,y)\mid Hg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),{H}\in\hat{\mathcal{C}}_{\mathrm{opt}}\}\vskip-2.84526pt$
@@
$\{(g,\sigma_{y},u,y)\mid\bar{H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),h(g)=0\},\vskip-2.84526pt$
where $\bar{H}$ is a fixed trajectory matrix. We next move the constraint on $g$ to the cost via regularization, leading to a direct DDPC
@@
$\mathrm{subject~to}$ $\bar{H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
@@
Similar to (13), to ensure the equivalence of the optimal solution, one may require the regularization parameter $\lambda_{w}$ in (16) to approach infinity. Fortunately, in our context, we can establish an exact penalty with a finite regularization parameter $\lambda_{w}$. In particular, consider a quadratic optimization problem of the form
@@
Consider (17) and (18), where $\mathcal{X}$ is a convex set and $M$ is positive semidefinite. Then, for any fixed $\lambda_{w}\geq 0$, (18) is a convex relaxation of (17), i.e., the optimal value of (18) is no larger than that of (17). Further, if there exists an optimal solution $(x_{1}^{*},x_{2}^{*})$ for (17) such that $x_{1}^{*}\in\mathcal{X}^{\circ}$, then there exists a finite $\lambda_{w}^{*}>0$, such that (17) and (18) are equivalent for all $\lambda_{w}>\lambda_{w}^{*}$, i.e., they have the same optimal solutions.
It is clear that (18) is a convex relaxation of (17) since every optimal solution of (17) is feasible for (18) with the same objective value. One key fact in this theorem is that we can find a finite penalty parameter such that (17) and (18) are equivalent. Theorem 1 confirms that $\|Dx_{2}\|_{p}$ is an exact penalty function for the quadratic problem (17). Our proof adapts the notion of partial calmness [ye1997exact]. Let $f(x)=x^{{\mathsf{T}}}Mx$. The key idea is to show that any feasible point $(x_{1},x_{2})$ of the penalized problem around a local region of $(x_{1}^{*},x_{2}^{*})$ can be “repaired” to a nearby feasible point $(\tilde{x}_{1},\tilde{x}_{2})$ satisfying $Dx_{2}=0$. Then, we illustrate the difference between $f(x_{1})$ and $f(\tilde{x}_{1})$ is controlled proportionally to $\|Dx_{2}\|_{p}$. Combining it with the fact that $f(\tilde{x}_{1})\geq f(x_{1}^{*})$ yields a local lower bound of the form $f(x_{1})+\lambda_{w}\|Dx_{2}\|_{p}\geq f(x_{1}^{*})+(\lambda_{w}-\lambda_{w}^{*})\|Dx_{2}\|_{p}$, implying that for any $\lambda_{w}>\lambda_{w}^{*}$ it is never optimal to keep a nonzero constraint violation. Finally, the convexity of (18) promotes this local argument to a global equivalence. We present the full proof in Appendix A.
To illustrate Theorem 1, we randomly generate problem instances of (17) and (18). As shown in Figure 3, all $p$-norms render $\|Dx\|_{p}$ an exact penalty. Note that smaller values of $p$ lead to faster convergence of the optimal value of (18) toward that of (17) as $\lambda_{w}$ increases. Since $\|Dx\|_{p}$ decreases monotonically with $p\in[1,\infty]$, larger $p$ values impose weaker penalties and therefore require larger $\lambda_{w}$ to guarantee equivalence between the two problems.
@@ 4 Regularization in direct data-driven control via convex relaxations
In this section, we apply Theorem 1 and derive three direct DDPC strategies via convex relaxations, including regularization using projection-based norm, causality-based norm, and $l_{1}$ norm. The projection-based norm and $l_{1}$ norm are widely used in the literature (see [dorfler2022bridging] and its references), and the causality-based norm is new.
@@ 4.1 Projection-based norm and a Hankel-form SPC
which only considers the SPC constraint (10b) for the inner identification. This bi-level DDPC (19) is much simpler than the general one (7). However, for a fixed $\lambda_{y}>0$, the optimal value of (19) is not necessarily lower than that of (7). This is because dropping constraints in the inner problem (10) does not necessarily enlarge its optimal solution set, which constitutes the feasible region of the outer problem. Therefore, similar to Example 1, (19) is not a relaxation of (7).
@@
$\Pi_{1}:=H_{1}^{\dagger}H_{1},\qquad H_{1}:=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}).\vskip-2.84526pt$ (20)
@@
Note that (21) differs slightly from the classical SPC formulation, and we call it Equality-form SPC. In particular, the classical SPC does not involve the use of an identified trajectory library $H_{\textnormal{s}}^{*}$; instead, it employs an explicit multi-step predictor (also see Remark 2). We provide further details in Appendix B.1.
Following the discussion in Section 3.3, we here fix $H=H_{\mathrm{d}}$ and impose a constraint on $g$ to get an equivalent reformulation for (21). In this case, we use a projection-based norm $\|(I-\Pi_{1})g\|_{2}$ as in [dorfler2022bridging], and consider
@@
The proof follows a similar idea in [shang2024convex], and we postpone it to Appendix B.2. The key idea is to show that feasible regions for decision variables $u,y,\sigma_{y}$ are the same for (19), (21), and (22). Note that the data matrix $H_{\textnormal{d}}$ in Proposition 1 can be arbitrary and may come from a nonlinear system.
It is now clear that (22) is in the form of (17) (we give a detailed construction in Appendix B.4). Thus, applying Theorem 1 leads to the following result.
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},$ and $\mathcal{Y}$ are convex sets. Consider the regularized problem
@@
For any $\lambda_{g}\geq 0$, it is a convex relaxation of (22). If the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ to (22) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$ such that (22) and (23) are equivalent for all $\lambda_{g}>\lambda_{g}^{*}$. In other words, they have the same optimal cost value and optimal solution.
We call (23) Linearity-based DDPC (L-DDPC). Combining Corollary 1 with Proposition 1 leads to the result below.
@@
Consider the bi-level DDPC (19) and the single-level L-DDPC (23). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are convex sets. Then, for any nonnegative $\lambda_{g}\geq 0$, (23) is a convex relaxation of (19), that is:
@@
Our Theorem 2 is closely related to [dorfler2022bridging, Theorem 4.6], but there are two key differences. Conceptually, as illustrated in Example 1, the L-DDPC (23) is only a convex relaxation of (19), and not necessarily of the original bi-level problem (7). This subtle distinction was overlooked in [dorfler2022bridging, Theorem 4.6]. Technically, we establish that (23) serves as a convex relaxation of (19) for any nonnegative parameter $\lambda_{g}\geq 0$, rather than only for sufficiently small $\lambda_{g}$ as in [dorfler2022bridging, Theorem 4.6]. The threshold $\lambda_{g}^{*}$ that makes (19) and (23) equivalent depends on both the Lipschitz constant of the cost function over a specific region and problem data (e.g., $H_{\textnormal{d}}$) rather than only on the Lipschitz constant as in [dorfler2022bridging, Theorem 4.6]. Moreover, our proof is elementary, relying on Theorem 1 and Proposition 1, which are inspired by the notion of partial calmness. We avoid utilizing [clarke1990optimization, Proposition 2.4.3] as in [dorfler2022bridging, Theorem 4.6], whose application in the DDPC setting may require additional adaptations.
@@ 4.2 Causality-based norm and a Causal SPC
which considers the SPC constraint (10b) and the causality constraint (10c) for the inner identification. If $H_{\textnormal{d}}$ has full row rank, the optimal solution $H^{*}_{\textnormal{sc}}$ for the inner problem in (24) is unique [sader2023causality, Lemma 2], which can be obtained via its optimal solution $K^{*}_{\textnormal{sc}}$, derived as
$K_{\textnormal{sc}}^{*}=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1},\vskip-2.84526pt$
@@
$H_{\textnormal{d}}=\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}\\ L_{31}&L_{32}&L_{33}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\end{bmatrix}\vskip-2.84526pt$
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are convex sets. Fix any data matrix $H_{\textnormal{d}}$ in (5b) with full row rank and sufficiently large column number. The bi-level problem (24) has the same optimal value as the single-level problems (26) and (28). Furthermore, all three problems have the same unique optimal solution $u^{*},y^{*}$, and $\sigma_{y}^{*}$.
The idea is to show (24), (26) and (28) share the same feasible region; see Appendix B.3. It is clear that (28) is in the form of (17). Thus, applying Theorem 1 leads to the following result.
@@
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U}$, and $\mathcal{Y}$ are convex sets. Consider the regularized problem
@@
with $\hat{H}$ given in (27). For any $\lambda_{g}\geq 0$, it is a convex relaxation of (28). Further, if the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (28) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$, such that, problem (29) has the same optimal value and optimal solutions as (28) for all $\lambda_{g}>\lambda_{g}^{*}$.
@@
Consider the bi-level DDPC (24) and the single-level C-DDPC (29). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $H_{\textnormal{d}}$ has full row rank with sufficiently large column number. Let $\Gamma,\mathcal{U}$ and $\mathcal{Y}$ be convex sets. For any nonnegative $\lambda_{g}\geq 0$, (29) is a convex relaxation of (24), that is:
@@
Theorems 2 and 3 show the direct L-DDPC and C-DDPC are convex relaxations of the corresponding bi-level DDPC formulations. Since hard identification constraints are replaced by regularization terms, these direct approaches admit flexibility in terms of soft identification. Corollaries 1 and 2 developed from Theorem 1 provide theoretical conditions that characterize when the relaxed problems become equivalent to the original formulations.
Remark 3 (Comparison with [sader2023causality]).
To our knowledge, the single-level C-DDPC in (29) is new, and its relationship with the bi-level DDPC (24) has not been studied previously. The closest related work is [sader2023causality], which introduces a causality-based regularizer within the $\gamma$-DDPC framework. The decision variable is reformulated and the regularization is imposed on $\gamma$, thereby losing the explicit trajectory library interpretation. Instead, our method (29) directly regularizes $g$, which retains the structure of the trajectory library. In particular, C-DDPC can then be naturally combined with the $\|g\|_{1}$ term, which implicitly performs trajectory selection, as we will discuss in Section 4.3.
@@ 4.3 Low-rank approximation and $l_{1}$ norm
$H_{\textnormal{lr}}^{*}=\sum_{i=1}^{mL+n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}\vskip-5.69054pt$
@@
$\mathrm{subject~to}$ ${H}_{\textnormal{lr}}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
@@
Proof.
Let $(g^{*},\sigma_{y}^{*},u^{*},y^{*})$ be an optimal solution of (31), and define $b^{*}:=H_{\textnormal{lr}}^{*}g^{*}.$ Since $\mathrm{Im}(H_{\textnormal{lr}}^{*})\subseteq\mathrm{Im}(H_{\textnormal{d}})$, there exists some $\hat{g}$ such that $H_{\textnormal{d}}\hat{g}=b^{*}.$ Hence $(\hat{g},\sigma_{y}^{*},u^{*},y^{*})$ is feasible for (32). Since the objective does not depend on $g$, it has the same objective value. Therefore, if $\alpha_{\textnormal{c}}\geq\|\hat{g}\|_{1}$, then $(\hat{g},\sigma_{y}^{*},u^{*},y^{*})$ is feasible for (32), and we thus complete the proof.
@@
Let $Q\succ 0,R\succ 0,\lambda_{y}>0$, and $\Gamma\times\mathcal{U}\times\mathcal{Y}$ being convex. Suppose problem (32) satisfies Slater’s condition. Consider the regularized problem
@@
There exists a penalty parameter $\lambda_{g}^{*}\geq 0$, such that any optimal solution $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ of (32) is also an optimal solution of (33) for $\lambda_{g}=\lambda_{g}^{*}$.
Proof.
To simplify notation, we let $x:=\textnormal{col}(\sigma_{y},u,y),\mathcal{C}:=\Gamma\times\mathcal{U}\times\mathcal{Y},$ and define $f(x):=\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}.$ The equality constraint in (32) can be written as $H_{\textnormal{d}}g-Ax=b$ for some matrix $A$ and vector $b$.
Then we rewrite (32) as
$\min_{x\in\mathcal{C},g}\ f(x)\quad\text{subject to}\quad H_{\textnormal{d}}g-Ax=b,\quad\|g\|_{1}\leq\alpha_{\textnormal{c}}.\vskip-5.69054pt$
Strong duality holds for (32) by Slater’s condition. Let $(\lambda_{g}^{*},\mu^{*})$ be a dual optimal solution, with $\lambda_{g}^{*}\geq 0$. Take any primal optimal solution $(x^{*},g^{*})$. By strong duality of (32), we know that $(x^{*},g^{*})$ minimizes the Lagrangian
$f(x)+\lambda_{g}^{*}(\|g\|_{1}-\alpha_{\textnormal{c}})+(\mu^{*})^{\top}(H_{\textnormal{d}}g-Ax-b).\vskip-5.69054pt$
Dropping the constant term $\lambda_{g}^{*}\alpha_{\textnormal{c}}$, we conclude that $(x^{*},g^{*})$ also minimizes $f(x)+\lambda_{g}^{*}\|g\|_{1}+(\mu^{*})^{\top}(H_{\textnormal{d}}g-Ax-b).$ Since $(x^{*},g^{*})$ is feasible for (33), it follows that $(x^{*},g^{*})$ is optimal for (33). As $(x^{*},g^{*})$ is an arbitrary primal optimal solution of (32), the result follows.
We note that the converse implication generally does not hold in Proposition 4: an optimal solution of (33) need not be feasible, and hence need not be optimal, for (32). Propositions 3 and 4 imply that (33) is a relaxation of (30) with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ under mild conditions.
@@
Consider the bi-level problem (30) and the single-level problem (33). Suppose $Q\succ 0$, $R\succ 0$, $\lambda_{y}>0$, and $\Gamma$, $\mathcal{U}$, and $\mathcal{Y}$ are convex sets. Let $\lambda_{g}^{*}\geq 0$ be the penalty parameter given by Proposition 4. Then, for any $0\leq\lambda_{g}\leq\lambda_{g}^{*}$, problem (33) is a convex relaxation of (30) with respect to the function $f(\sigma_{y},u,y):=\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2},$ in the following sense:
@@
    any feasible $(\sigma_{y},u,y)$ for (30) is feasible to (33);
@@
    for any optimal solution $(g^{*},\sigma_{y}^{*},u^{*},y^{*})$ of (33), $f(\sigma_{y}^{*},u^{*},y^{*})$ is smaller than the optimal value of (30).
A similar statement as Theorem 4 appears in [dorfler2022bridging, Theorem 4.8]. However, we emphasize that the relaxation is only true with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ under an upper bound for the regularizing parameter $\lambda_{g}$, which is treated differently in [dorfler2022bridging]. The reason is that while the optimal solution of (33) is the same as (32), their objective functions are different, which generally leads to different optimal values.